Determine whether f is a function from the integers to the set of all real numbers.
Enter "Y" for yes and "N" for no.
1. f(n)=±n
2. f(n)=1/(n^2−16)
3. f(n)=√n^2+6
4. f(n)=1/(n^2+9)
1
Expert's answer
2020-10-26T20:05:03-0400
  1. "f(n)=\pm n" is not a function, since each number corresponds to two different numbers.
  2. "f(n)=\frac{1}{\sqrt{n^2-16}}" is defined for all numbers, except 4 and therfore is not a function.
  3. "f(n)=\sqrt{n^2}+6" is defined for all numbers. It is a function.
  4. "f(n)=\frac{1}{\sqrt{n^2+9}}" is defined for all numbers. It is a function.